OpenStudy (anonymous):

im having trouble Find dy/dt. y = cos^6(piΔt - 9) A) -6 cos^5(pi* t - 9) sin(pi*t - 9) B) -6 picos^5(pi *t - 9) sin(pi*t - 9) C) 6 cos^5(pi *t - 9) D) -6 sin^5(pi *t - 9)

OpenStudy (anonymous):

Use the chain rule.

OpenStudy (anonymous):

take me through the steps?

OpenStudy (anonymous):

none of these look right to me, but maybe i don't understand the question

OpenStudy (anonymous):

uhh let me check

OpenStudy (anonymous):

The 14) should be on a separate line.

OpenStudy (anonymous):

what is the 14 doing out at the end?

OpenStudy (anonymous):

yeh it shouldnt be there sorry

OpenStudy (anonymous):

those triangle should be empty spaces

OpenStudy (anonymous):

There should be a factor of pi multiplying the answer.

OpenStudy (anonymous):

ok now what about the \(\pi t\) ? if it is there, there should be a coefficient including \(\pi\) out front somewhere

OpenStudy (anonymous):

in the question there isnt one

OpenStudy (anonymous):

Which is the problem we (answering it) are having.

OpenStudy (anonymous):

cos^6(pit-9)

OpenStudy (anonymous):

well what is dy/dt?

OpenStudy (anonymous):

\[f(x)=\cos^6(\pi t-9)\] ?

OpenStudy (anonymous):

\[y=\cos^6 (\pi t -9)\] \[ =[\cos(\pi t -9)]^6\]

OpenStudy (anonymous):

so what does that mean exactly?

OpenStudy (anonymous):

Ah nvm what I said before. You just use the chain rule.

OpenStudy (anonymous):

the answer that i have doesnt match any of hte available

OpenStudy (anonymous):

6 π cos^5(9-π t) sin(9-π t)

OpenStudy (anonymous):

In the parenthesis it should be pi t minus nine, not nine minus pi t.

OpenStudy (anonymous):

Yep there should be a pi in front of the 6.

OpenStudy (anonymous):

so im clueless

OpenStudy (anonymous):

A) -6 cos^5(pi* t - 9) sin(pi*t - 9) B) -6 cos^5(pi *t - 9) sin(pi*t - 9) Both A and B are the same.

OpenStudy (anonymous):

You must of done a typo youself or if you copied correctly, one of them should have a pi in front of the -6.

OpenStudy (anonymous):

yeh i see one has a pi after the six and a doesnt

OpenStudy (anonymous):

A) -6 cos^5(pi* t - 9) sin(pi*t - 9) B) -6 picos^5(pi *t - 9) sin(pi*t - 9)

OpenStudy (anonymous):

Now that's better.

OpenStudy (anonymous):

Could you show me you working now?

OpenStudy (anonymous):

So you use the chain rule and what did you get?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

Remember: \[\frac{d}{dx}\cos x =-\sin x\]

OpenStudy (anonymous):

That's for reference to when you're differentiating cosine.

OpenStudy (anonymous):

d/dt(cos^6(pit-9)) =d/dt(cos^6(9-pit)) =6 cos^5(9-pit)[d/dt(cos(9-pit))] =-d/dt(9-pit)sin(9-pit) 6cos^5(9-pit) =-6 sin(9-pit) cos^5(9-pit)[d/dt(9-pit)] =d/dt(9)-pid/dt(t) -6 cos^5(9-pit) sin (9-pit) =-6 sin(9-pit) cos^5(9-pit) [0-pi[d/dt(t)]] =6 pi sin(9-piy) cos^5(9-pit)[d/dt(t)] =1 6 pi cos^5(9-pit) sin(9-pit) =6 pi sin(9-pit)cos^5(9-pit)

OpenStudy (anonymous):

Why'd you change the pit-9 to 9-pit? Seems like a habit for you.

OpenStudy (anonymous):

You don't need ot change it around and do unnecessary switcheroos.

OpenStudy (anonymous):

need to*

OpenStudy (anonymous):

ill be honest that is what the computation calculator does

OpenStudy (anonymous):

Computation calculator. Did you do it yourself using pen and paper? It's much easier for you to do this with pen and paper.

OpenStudy (anonymous):

i used the calculator because i dont know how to do it

OpenStudy (anonymous):

Okay, let's see: \[y = \cos^6(\pi t - 9)\] \[y=[\cos(\pi t -9)]^6\] Have you learnt the chain rule yet? If you have, then have you paid attention in class about moving the power to the front and differentiating the inside, etc.?

OpenStudy (anonymous):

no i did not pay attention

OpenStudy (anonymous):

Is this familiar with you or not? Because I need to know whether you've done questions that involve differentiation with the chain rule.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

This is tricky, because if you fall behind in maths, you will struggle. Okay... Can you differentiate this using the chain rule? \[\frac{d}{dx}(x-1)^2=?\]

OpenStudy (anonymous):

maybe the calculator can

OpenStudy (anonymous):

No, please don't use the calculator...

OpenStudy (anonymous):

Would you be allowed to have a calculator of this kind when you do a test/exam?

OpenStudy (anonymous):

The answer would most likely be NO. So for your benefit, you need to do this with pen and paper.

OpenStudy (anonymous):

this is one question out of like 28 for my final which is due tomorrow by midnight. i had 6 chapters to read this semester and only read 3. it is an online class so i submit tomorrow night

OpenStudy (anonymous):

i never finished chain rule i dont think but it is too late now

OpenStudy (anonymous):

Did procrastination lead to this? SHould of started when you got the final.

OpenStudy (anonymous):

i have to try to pass this final with as much correct answers. I KNOW i got it on like last weekend but waited until like today or yesterday to start

OpenStudy (anonymous):

Don't teachers teach you? I find many students on OpenStudy learn online.

OpenStudy (anonymous):

We're not responsible for your wrongdoing. We're only here to help guide you with questions.

OpenStudy (anonymous):

That's why people with that attitude won't do well. You have to ASK the TEACHER for help and assistance when unsure of anything. Even if it was simple knowledge. the teacher gets paid for all this.

OpenStudy (anonymous):

The chain rule is simple if you can follow this: \[\frac{d}{dx}(x-1)^2=2(x-1)^{2-1}\times ?\] First I rewrote the expression. Then I moved the power to the front. Then I subtracted 1 from the power. Now you can see there is a question mark which indicates somethings missing there when I'm differentiating using chain rule. I will talk about this in the next step. Okay: \[\frac{d}{dx}(x-1)^2=2(x-1)^{2-1}\times \frac{d}{dx}(x-1)\] As you can see, I filled in the question mark with the derivative of x-1. That is simply what's inside (x-1)^2. So basically that is all you need to know for the chain rule. Here's a summary of the steps in using the chain rule. 1. Rewrite expression 2. Bring the power to the front. 3. Subtract one from the original power. 4. Multiply that result with the derivative of what's inside the paranthesis/brackets.

OpenStudy (anonymous):

Apply this to your question as well as using your knowledge of differentiating trig functions.